### v2

# Large Language Models Badly Generalize across Option Length, Problem Types, and Irrelevant Noun Replacements

Abstract
In this paper, we propose a “Generalization Stress Test” to assess Large Language Models’ (LLMs) generalization ability under slight and controlled perturbations, including option length, problem types, and irrelevant noun replacements. We achieve novel and significant findings that, despite high benchmark scores, LLMs exhibit severe accuracy drops and unexpected biases (e.g., preference for longer distractors) when faced with these minor but content-preserving modifications. For example, Qwen 2.5 1.5B’s MMLU score rises from 60 to 89 and drops from 89 to 36 when option lengths are changed without altering the question. Even GPT4o experiences a 25-point accuracy loss when problem types are changed, with a 6-point drop across all three modification categories. These analyses suggest that LLMs rely heavily on superficial cues rather than forming robust, abstract representations that generalize across formats, lexical variations, and irrelevant content shifts.
Large Language Models Badly Generalize across Option Length, Problem Types, and Irrelevant Noun Replacements
1 Introduction
Large Language Models (LLMs) have achieved near-human performance across a variety of natural language processing (NLP) benchmarks, from elementary tests (Cobbe et al., 2021) to university-level challenges (Hendrycks et al., 2021). This success has spurred claims that LLMs are approaching human-like generalization capabilities (OpenAI, 2024; Bubeck et al., 2023; Jones and Bergen, 2024). However, it remains unclear whether their high benchmark scores reflect genuine generalization or if LLMs are simply exploiting superficial cues that fail under slight perturbations.
While LLMs perform well in established benchmarks, concerns have been raised about the validity of these evaluations Chen et al. (2023); Ye et al. (2023). Data contamination, where models unintentionally learn from benchmark data included in their training, can inflate performance estimates (Brown et al., 2020; Xu et al., 2024; Ravaut et al., 2024; Zhou et al., 2023). These issues suggest that existing benchmarks have exposed patterns and may not truly assess generalization.
Recent work has focused on uncovering the actual limits of LLM generalization. One direction involves the development of dynamic evaluation methods that modify the evaluation process on the fly (Zhu et al., 2024; Yu et al., 2024). Another approach emphasizes creating more challenging or adversarial test sets that push models beyond their current capabilities, such as MMLU-Pro (Wang et al., 2024) and GSM-Plus (Li et al., 2024a). A third line of inquiry involves introducing subtle modifications to benchmark datasets to test LLM robustness, such as altering the order of multiple-choice options or changing the format of questions (Zheng et al., 2024; Li et al., 2024b; Gupta et al., 2024; Alzahrani et al., 2024). While these approaches have contributed to a better understanding of LLM performance, they either totally change the original problems, increase the complexity of the evaluation, or focus on relatively limited formatting changes like option ID adjustments.
Figure 1: Generalization stress tests and summarized results. LLMs do not generalize well across various option lengths, problem types, and noun replacements. Tested models are Qwen2.5 1.5B, 7B, 72B, and GPT4o.
We find serious biases of recent SoTa LLMs to common patterns by introducing an evaluation framework, Generalization Stress Tests, which examines LLMs under three types of minor, content-preserving perturbations:
  • Altering option length (e.g., increasing the length of distractors or correct options without changing their semantic content).
  • Changing problem types (e.g., converting multiple-choice questions to boolean questions).
  • Replacing irrelevant nouns (e.g., substituting semantically irrelevant nouns in prompts).
As shown in Figure 1, these simple modifications, surprisingly, lead to substantial performance degradation [^1] We test GSM-8K for noun replacement, as some MMLU cases lack irrelevant nouns.. We observe that LLMs struggle to generalize across varying option lengths, problem types, and noun replacements. For example, Qwen 2.5 1.5B’s MMLU score drops from 89 to 36 when option lengths are changed without altering the question. Even GPT4o experiences a 25-point accuracy loss when question types are changed, with a 6-point drop across all three categories. These findings reveal a critical limitation: LLMs are biased to specific irrelevant patterns and fail to replicate the human-like ability to ignore irrelevant format details.
2 Methods: Generalization Stress Tests
We conduct generalization stress tests by applying minor modifications to the original benchmark, focusing on variations in option length, scoring type, and the replacement of irrelevant nouns.
We investigate typical tasks for LLMs that include multiple-choice questions (MCQ) and open-ended question answering (Open-ended QA).
2.1 Alter Option Length to Analyze LLMs’ Length Bias
Make the right option longer (RL): Question: What is the capital of France? A) Berlin B) Madrid C) Paris, a city renowned for its art, fashion, and cuisine. D) Rome Make one wrong option longer (WL): Question: What is the capital of France? A) Berlin, known for its vibrant culture and historical landmarks. B) Madrid C) Paris D) Rome
Figure 2: An illustration of altering option length. The ground truth of this question is C) Paris.
To analyze whether LLMs are generalized across option length or whether LLMs are biased toward long options in MCQ. We first make all options in a problem longer by asking GPT4o [^2] We use its API version provided by Microsoft Azure. to make the options longer without including information that could help answer the question. Refer to Appendix A for generation details.
As illustrated in Figure 2, we then design the following two types of lengthening problems: a)Make one wrong option longer (WL), b)Make the right options longer (RL).
Length Control: To assess the impact of option length on LLM generalization, we control the length of the lengthened options in the WL condition. Specifically, we ask GPT4o to generate options of varying lengths: (a) < 10 tokens, (b) 10 to 20 tokens, and (c) > 20 tokens.
Paraphrase Verification: We also enlist human experts to verify whether the paraphrased options do not introduce unintended biases or hints. Details can be found in the Appendix A.
2.2 Change Problem Type to Fairly Analyze LLMs’ Scoring Bias
Cloze: Question: What is the capital of France? Answer: _ (Selected from whole vocabulary) Bool questions: 1. Question: What is the capital of France? Answer: Paris The answer is _(Selected from True/False) 2. Question: What is the capital of France? Answer: Berlin The answer is _(Selected from True/False) Require to judge both two propositions correctly.
Figure 3: An illustration of changing the scoring type from MCQ to bool questions.
Previous work found LLMs do not generalize to different option IDs in MCQ Zheng et al. (2024) and tried to solve this by changing the task to cloze Alzahrani et al. (2024). However, the cloze task reduces the expected value of selecting the correct answer. Therefore, we propose changing the multiple-choice questions to Boolean questions, requiring two judgments to be accurate, so that the difficulty of the questions is as similar as possible to that of multiple-choice questions.
As illustrated in Figure 3, we derive one true proposition that concludes with the right option and one false proposition that is a randomly selected wrong option.
2.3 Replace Irrelevant Nouns to Analyze Bias towards Irrelevant Content
Problem with irrelevant noun: Question: John lives in France; what is his country’s capital? A) Berlin B) Madrid C) Paris D) Rome Answer: C Problem after modifying the irrelevant noun: Question: Mike lives in France; what is his country’s capital? A) Berlin B) Madrid C) Paris D) Rome Answer: C
Figure 4: An illustration of replacing irrelevant nouns.
In open-ended QA like those in GSM8K Cobbe et al. (2021), the questions may contain nouns that are unrelated to the answers. In this subsection, we explore the impact of changes to these unrelated nouns on the decision-making of large models. As shown in Figure 4, we replaced nouns in the questions, such as names of people and animals, ensuring that these replacements do not alter human decision-making. Details are in Appendix B.
Semantic relevance control Additionally, regarding noun replacements, we also examined the impact of the semantic proximity of the replacements. We conducted experiments in this area by instructing GPT-4o mini to perform replacements with varying degrees of semantic similarity.
3 Experiments
We perform evaluations on harness framework Gao et al. (2024) and adopt its default setting. We evaluate models of Llama3.1 series Dubey et al. (2024), Qwen2.5 series Yang et al. (2024b), and GPT4o. Llama3.1, and Qwen2.5 are the most powerful small models, while GPT4o is the most powerful LLM. We evaluate LLMs on MMLU Hendrycks et al. (2021), ARC-Challenge Clark et al. (2018), and GSM8k Cobbe et al. (2021). The first two are MCQ benchmarks, and the last consists of open-ended QAs. Refer to Appendix C for detailed experimental setups.
3.1 Results of Altering Option Length
Benchmark Model Origin RL WL
MMLU Qwen2.5 1.5B 60.3 89.0 36.3
Qwen2.5 7B 73.7 90.1 55.6
Qwen2.5 72B 85.4 94.1 75.6
LLaMa3.1 8B 65.5 85.6 53.6
LLaMa3.1 70B 78.8 93.6 70.6
GPT4o mini 76.5 87.2 70.6
GPT4o 85.2 89.7 83.3
ARC-C Qwen2.5 1.5B 77.3 88.9 68.1
Qwen2.5 7B 90.0 94.3 84.0
Qwen2.5 72B 95.8 97.2 94.4
LLaMa3.1 8B 78.1 85.2 74.7
LLaMa3.1 70B 91.8 96.3 90.8
GPT4o mini 91.8 95.1 91.4
GPT4o 96.5 97.1 95.5
Table 1: Performance on altering option length. RL refers to lengthening the right option; WL refers to lengthening the wrong option. The values are percentages.
LLMs struggle to generalize across option length: From Table 1, it is evident that across all LLMs, from 1.5B to GPT4o, scores increase significantly when the length of the correct option is extended and decrease significantly when we make an incorrect option longer. Smaller models generalize even worse. In Appendix D.1, we introduce another setting of making all options longer, in which our finding that LLMs are biased towards the longer option persists.
Length matters, especially when we lengthen the right option. As shown in Table 2, changing the length can result in a difference of more than 10 points in the RL setting.
Settings <10 10 to 20 >20
Origin 65.5%
RL 70.0% 75.3% 84.0%
WL 64.5% 60.7% 61.6%
Table 2: The performance of LLaMa3.1 8B on MMLU changes when gradually altering the length of correct and incorrect options.
Another intriguing finding is that LLMs tend to select the right option if we make all incorrect options longer, refer to Appendix D.2.
3.2 Results of Altering Scoring Type
LLMs do not have invariant knowledge that can generalize across scoring types. As in Table 3, all models tend to score lower when the benchmarks are changed from the original format to boolean questions. Qwen2.5 1.5B and Llama3.1 8B score only half the points in the MMLU’s "both" setting. Smaller models generalize worse.
Benchmark Model MCQ BQ Both
MMLU Qwen2.5 1.5B 58.8 30.3 22.1
Qwen2.5 7B 72.4 54.7 46.7
Qwen2.5 72B 84.0 69.1 65.0
LLaMa3.1 8B 64.6 40.6 32.6
LLaMa3.1 70B 78.4 63.5 56.7
GPT4o mini 75.1 54.5 49.2
GPT4o 84.7 59.5 56.8
ARC-C Qwen2.5 1.5B 74.0 40.4 35.2
Qwen2.5 7B 89.5 69.4 66.4
Qwen2.5 72B 95.0 85.8 84.4
LLaMa3.1 8B 77.4 53.6 47.1
LLaMa3.1 70B 92.1 82.7 79.2
GPT4o mini 90.6 79.7 76.6
GPT4o 96.2 79.6 76.2
Table 3: Performance on changing problem type from multi-choice question (MCQ) to bool questions (BQ). The values are percentages.
3.3 Results of Replacing Irrelevant Nouns
Models Origin Replace Nouns
Qwen2.5 1.5B 62.5% 54.9%
Qwen2.5 7B 83.5% 78.0%
Qwen2.5 72B 92.3% 81.9%
Llama3.1 8B 54.7% 51.7%
Llama3.1 70B 80.8% 74.2%
GPT4o mini 71.3% 64.1%
GPT4o 86.7% 79.5%
Table 4: Performance of replacing nouns on GSM8K. We report results on it since it has irrelevant nouns.
Replacing irrelevant nouns degrades performance consistently across various models. As seen in Table 5, the scores of all models drop when the terms are renamed, with the magnitude of the decrease being similar across models. GPT4o models still show a decline.
Models Origin High Medium Low
Llama3.1 8B 54.7% 51.5% 48.0% 44.0%
Qwen2.5 7B 83.5% 82.0% 78.1% 70.7%
Table 5: Model performance on replacing nouns with various semantic relevance levels.
Replacing irrelevant nouns with semantically distant words further reduces the effectiveness.
3.4 Reasoning Behind Accuracy Drops
The above analysis reveals that LLMs are severely biased to common but irrelevant patterns.
4 Conclusion
This paper finds that LLMs exhibit significant performance degradation when faced with slight changes in question format, option length, or irrelevant content shifts. These findings underscore that LLMs rely on superficial patterns rather than robust, generalizable reasoning. By introducing the "Generalization Stress Tests," we offer novel ideas towards evaluating LLMs’ true generalization capabilities.
Limitations
This work focuses solely on non-chain-of-thought LLMs, such as GPT-4o, and does not consider emerging O1.
Ethnic Statement
This work adheres to ACL’s ethical guidelines, we state that there are no ethical concerns to our knowledge.
Appendix A Prompts and Verification in Altering Option Length
A.1 Prompts
We chose the GPT-4o to lengthen options.
The default prompt to lengthen options is: The user will give you a question, the choices, and the answer from a dataset. Rewrite the four choices into longer ones. Make sure not to change the question willingly. Make sure that the rewritten options do not contain a hint of the correct answer.
The prompt to control option length is: We concatenate the default prompt to one of the following prompts.
  • Make sure that each rewritten option contains no more than 10 words.
  • Make sure that each rewritten option at least 10 words and no more than 20 words.
  • Make sure that each rewritten option contains at least 20 words.
We set the temperature to 0, and the other setting is the same as the default.
A.2 Verification Process
We manually verified the rewritten sentences to check whether lengthening the sentence introduced factors related to the answer or changed the question’s meaning. We manually checked 100 examples from MMLU and found that 99 had no issues, while 1 changed the original meaning of the question. The rewriting accuracy was 99%.
Appendix B Prompts in Replacing Irrelevant Nouns
We found that GPT-4o and GPT-4o mini perform similarly on this task. To reduce carbon emissions, we chose the GPT-4o mini.
The prompt to simply replace irrelevant nouns is: Assist in creatively substituting nouns in mathematical problems to prevent students from memorizing solutions. The replacements should be imaginative, ensuring the mathematical relationships and the accuracy of the solutions are preserved. “input_text” Other than replacing nouns, do not alter the original word order sentence structure, or add or remove any sentences. Give the modified question directly.
The prompt to alter semantic relevance is: Substitute nouns and some relevant words in the mathematical problems creatively to prevent students from memorizing solutions. The replacements should be done in three levels:
  • Level 1: Only replace nouns with semantically similar words (e.g., ’apple’ becomes ’banana’).
  • Level 2: Replace nouns and verbs with words that differ in meaning but are still within the realm of common sense (e.g., ’apple’ becomes ’elephant’, ’eat fruit’ becomes ’drink coke’).
  • Level 3: Replace words as much as possible with highly imaginative and fantastical words, if you think it still makes sense in mathematical problems. (e.g., ’apple’ becomes ’alien gemstone’).
Apart from replacing nouns and some relevant words, maintain the original word order, sentence structure, and do not add or remove any sentences. Give three modified sentences directly, one for each level, only separated by ’###’. Don’t return anything else including ’Level 1’, ’Level 2’, ’Level 3’ but only "###". This is the original question: input_text
We set temperature to $0.1$, top-p to $1$, top-k to 0, and repetition_penalty to 0.
Appendix C Experiment Setup Details
This section describes the foundational setup of our experiments and analyses, including the evaluation framework and methods we used and the benchmarks and models we evaluated.
C.1 Evaluation Protocol
We perform evaluations on harness framework Gao et al. (2024). We chose harness because it is a flexible, configurable, reproducible framework. Unless otherwise specified, our evaluations are conducted in a 5-shot manner, with few-shot examples drawn from the benchmarks’ corresponding training sets.
C.2 Models
We evaluate models of Llama3.1 series Dubey et al. (2024), Qwen2 series Yang et al. (2024a), and GPT4o. Llama3.1 and Qwen2.5 are the most powerful small models, while GPT4o is the most powerful LLM. We list all models below.
  • Llama3.1 8B, Llama3.1 70B;
  • Qwen2.5 1.5B, Qwen2.5 7B, Qwen2.5 72B;
  • GPT4o, GPT4o mini.
C.3 Benchmarks
We evaluate LLMs on MMLU, ARC, Helaswag, GSM-MCQ, and GSM8k. The first four are MCQ benchmarks, and the last consists of open-ended questions.
  • MMLU Hendrycks et al. (2021) is a multi-task benchmark that covers 57 tasks ranging from elementary to college level. These tasks cover multiple disciplines, e.g., math, physics, law, history, etc. The whole test set consists of 14,042 examples. Following common practice, we calculate the accuracy of each task and report the average score across all tasks.
  • ARC Clark et al. (2018) is also a multitask dataset that includes data from eight types of tasks, testing aspects such as common sense, multi-hop reasoning, and algebraic operations, with 3,548 samples. ARC has two subsets: one is ARC-Challenge (abbreviated as ARC-C), and the other is ARC-Easy (abbreviated as ARC-E). The challenge set includes only those data that cannot be answered through retrieval and word co-occurrence methods, making it more difficult.
  • GSM-8K Cobbe et al. (2021) examines multi-step math word problems, which are relatively easy and designed to be solvable by middle school students. GSM8K is presented in an open-ended question format, unlike multiple-choice questions. It consists of 1,319 test questions.
    C.4 Budget
    We performed experiments with an H800 GPU; the total cost of the experiments was about 1000 GPU hours.
Appendix D Additional Results
D.1 Making All Options longer
Benchmark Model Origin AL RL WL
MMLU Qwen2.5 1.5B 60.3 54.7 89.0 36.3
Qwen2.5 7B 73.7 69.2 90.1 55.6
Qwen2.5 72B 85.4 81.3 94.1 75.6
LLaMa3.1 8B 65.5 64.3 85.6 53.6
LLaMa3.1 70B 78.8 76.0 93.6 70.6
ARC-C Qwen2.5 1.5B 77.3 67.3 88.9 68.1
Qwen2.5 7B 90.0 85.3 94.3 84.0
Qwen2.5 72B 95.8 93.1 97.2 94.4
LLaMa3.1 8B 78.1 78.6 85.2 74.7
LLaMa3.1 70B 91.8 89.9 96.3 90.8
Table 6: Performance on altering option length. AL refers to lengthening all options. RL refers to lengthening the right option. WL refers to lengthening the wrong option. The values are percentages.
We can see from Table 6 that LLaMa is more robust than Qwen, and larger models are more robust than smaller models, when we make all options longer. Besides, even if we introduce the setting of AL, our conclusion that LLMs are vulnerable to option lengths and biased to long options is not changed.
D.2 Make All Wrong Options Longer
Model origin WL WL-ALL
Llama3.1 8B 65.5% 53.6% 64.8%
Llama3.1 70B 78.8% 70.6% 82.4%
gpt-4o 85.2% 83.3% 85.6%
Table 7: Results of making all wrong options longer on the MMLU benchmark.
Making all wrong options could expose the right answer. From Table 7, we can see that if all the incorrect options are lengthened, the model will choose the only correct option that hasn’t been lengthened.